OpenStudy (anonymous):

Use the shortcut rules to mentally calculate the derivative of the given function. f(x) = −x + 9/x + 1 Im not sure how to do 9/x, mainly

OpenStudy (anonymous):

\(\frac 9 x \) can also be written as \(9x^{-1}\).

OpenStudy (anonymous):

@bloopman So -1-9? which is f1(x)= -10?

OpenStudy (anonymous):

$$\frac{d}{dx}[9x^{-1}] = -9x^{-2}$$

OpenStudy (anonymous):

@bloopman oh yes so its -1-9x^-2?

OpenStudy (anonymous):

Yes.